OpenStudy (anonymous):

Solve for b. 8b = 48 A. 384 B. 56 C. 40 D. 6

OpenStudy (cookiemonster2412):

divide both sides by 8

OpenStudy (cookiemonster2412):

so 48/8=6

OpenStudy (cookiemonster2412):

D

OpenStudy (anonymous):

d

OpenStudy (anonymous):

thanks can you help with another

OpenStudy (cookiemonster2412):

yea

OpenStudy (anonymous):

Which equation can be used to solve the problem? How many 12-packs of juice boxes contain a total of 84 juice boxes? A. 12 + b = 84 B. 12 ÷ b = 84 C. 12b = 84 D. b ÷ 12 = 84

OpenStudy (anonymous):

can someone help me answer a question?

OpenStudy (anonymous):

the ????? are dividing

OpenStudy (cookiemonster2412):

C

OpenStudy (anonymous):

one more

OpenStudy (cookiemonster2412):

k

OpenStudy (anonymous):

Which measurement is the width of this rectangle? Area = 102 cm² A. 6 cm B. 8 cm C. 9 cm D. 12 cm

OpenStudy (anonymous):



OpenStudy (cookiemonster2412):

uwant me to find area

OpenStudy (anonymous):

no width

OpenStudy (cookiemonster2412):

k

OpenStudy (anonymous):

area is 102cm

OpenStudy (cookiemonster2412):

A

OpenStudy (anonymous):

thank you

OpenStudy (cookiemonster2412):

yw